Lisa will make punch that is 25% fruit juice by
adding pure fruit juice to a 2-liter mixture that
is 10% pure fruit juice. How many liters of pure
fruit juice does she need to add?

Answer by alicealc(293)

pure fruit juice means 100% fruit juice
let the pure fruit juice = x liter

100%*x + 10%*2 = 25%*(x + 2)
(omit all the % sign)
100*x + 10*2 = 25*(x + 2)
100x + 20 = 25x + 50
100x - 25x = 50 - 20
75x = 30
x = 30/75 = 10/25 = 2/5 = 0.4

so, she needs to add 0.4 liter of pure fruit juice